OpenStudy (anonymous):

where is f(x)=ln(7+x^2) concave up and concave down and what are the inflection points

OpenStudy (anonymous):

take the first and second derivative... what do you get?

OpenStudy (anonymous):

f'(x)= 2x/ (x^2-7) and f"(x)= -2(x^2-7)/ (x^2+7)^2

OpenStudy (merchandize):

Concavity is determined by the second derivative. Concave up = positive second derivative Concave down = negative second derivative. Note that f'(x)=(2x)/(7+x^2) f "(x)=(2(7+x^2)-2x(2x) )/ (7+x^2)^2 = (14-2x^2)/(7+x^2)^2 Now you can continue and solve the inequality f '(x)>0 and f "(x)<0.

OpenStudy (anonymous):

but then what is the interval that it is concave up?

OpenStudy (anonymous):

-2x^2+14>0, then x<+-sqrt7?

OpenStudy (anonymous):

notice f is defined for all reals so you don't have to worry about the domain of f.

OpenStudy (anonymous):

potential inflection points are as -sqrt7 and sqrt7 break up your interval using those points. pick a #, c, from those intervals and stick it into the second derivative. if f''(c) > 0 , concave up... cc down otherwise.